# Universal Framework to Evaluate Automotive Perception Sensor Impact on Perception Functions

### v1

@@ V-A Model Selection
V-B Inputs and outputs
Inputs to the model are the track histories of the target vehicle and its surrounding vehicles, taken for the time frames from $t=t-t_{h}$ to $t=t$ and consist of the $x,\ y$ co-ordinates as detailed in [4]. For our analysis, the track histories are derived in two ways.
@@ V-B Inputs and outputs
Positional co-ordinates from both approaches are converted to the target vehicle’s frame of reference before using as inputs to the model. The output of the model infers the conditional probability distribution $P\left(Y\middle|X\right)$. Therefore, the inferred future $x,\ y$ positions at each future predicted time-frame are the parameters of a bivariate Gaussian probability distribution, where $\mu^{t}$ is the mean vector and $\Sigma^{t}$ is the covariance matrix.
@@
V-C Scenario Generation
Three scenario categories are defined with two representing a road segment from the US Highway 101 and the remainder, a straight highway road segment. The scenarios cover the most influential positions for a lane change by a target vehicle to impact the trajectory of the ego car, i.e. a lane change into the same lane as the ego car, directly in front of it, and in front of the lead vehicle to the ego car.
By considering each scenario category for both directions, six driving scenarios are created with a target vehicle performing a lane change manoeuvre. The ego car is represented by a typical passenger vehicle type along with several different vehicle types representing the surrounding vehicles. The ego car is equipped with a radar sensor and a stereo camera sensor.
  • Scenario 1: Lane change to the Right - In front of the ego car (US101)
  • Scenario 2: Lane change to the Left - In front of the ego car (US101)
  • Scenario 3: Lane change to the Right - In front of the lead car (US101)
  • Scenario 4: Lane change to the Left - In front of the lead car (US101)
  • Scenario 5: Lane change to the Right – In front of the lead car (Straight road)
  • Scenario 6: Lane change to the Left - In front of the lead car (Straight road)
The highway road segment consists of three lanes in each driving direction. Supplementary Data Tables A.1 and A.2 illustrate scenarios 1 and 2 with detailed descriptions in line with the definition of a scenario in [29].
@@ V-C Scenario Generation
High-fidelity sensor models for the radar and the stereo camera from CM are mounted on the ego vehicle positioned as shown in Fig. 4. The radar uses signal-to-noise ratio for detecting the surrounding objects with consideration given to the propagation and atmospheric losses whereas the camera uses a bounding box of the object for calculating the occlusions to determine whether an object is detected. The detection confidence is compared with a predefined threshold, which for this study is set to 0.6. Tables I and II list the default sensor specifications for the radar and the camera sensors except for the technical parameter being changed in the study, horizontal FOV (HFOV) and the range of the sensor, which is fixed at 100 m.
@@
VI Results and Analysis
A test run is established using a predefined scenario with one sensor modality and the sensor’s HFOV set as a variable. It is then simulated iteratively varying the assigned value to the variable, HFOV, as listed in Table III. Identical sets of simulation test runs are conducted for the radar and the camera sensors separately.
@@ VI-1 Impact of the Radar sensor’s ‘HFOV’ on Trajectory Prediction
Figs. 5, 6 and 7 show the derived RMSE values based on the radar sensor plotted against the prediction horizon of 5 s for the three scenario categories.
Figure 5: Impact on RMSE - Changing Radar’s HFOV: RLC Vs LLC - In front of Ego (US101)
Figure 6: Impact on RMSE - Changing Radar’s HFOV: RLC Vs LLC - In front of Lead (US101)
Figure 7: Impact on RMSE - Changing Radar’s HFOV: RLC Vs LLC - In front of Lead (Straight Road)
It is observed that the lowest setting of ${30}^{\circ}$ for the HFOV has produced the RMSEs closest to the ground truth-based RMSEs in all three scenario categories explored: (i) in front of the ego car on US101, (ii) in front of the lead car on US101 and (iii) in front of the lead car on the straight road. The intuitive expectation is for the sensor-based data obtained with a larger HFOV to perform better when used as the input to the prediction model. However, the results exemplify that the RMSEs of the predicted trajectories are optimal when the measurements are taken with the lowest HFOV of ${30}^{\circ}$.
It is believed since the dominant manoeuvre of all six scenarios is a lane change into the ego car’s lane, the narrow focus given by a smaller HFOV of the radar sensor enables the DL model to make a better prediction by reducing the complexity and unnecessary noise. Since for a lane change, the highest impact on the target vehicle is imposed by the vehicles in its closest proximity and directly ahead, a narrow HFOV improves focus by reducing the irrelevant data of unrelated surrounding traffic. Another likely rationale is that the data captured at ${30}^{\circ}$ HFOV is a close representative of the relevant area of interest around the target vehicle for the prediction model. The CS-LSTM model uses the BEV representation of the scene with a predefined boundary around the target car. Therefore, when the sensor data is based on a wider HFOV, the inputs may appear distorted reducing the model’s performance.
The above observations highlight the significance of determining the value addition by deploying high-specification perception sensors within the sensor suite of an AAD vehicle as mentioned in [17]. Although the expectation is that high-specification sensors would lead to high-quality data, the results of this experiment indicate that the performance of DL models might not always positively correlate with data from high-specification sensors.
@@ VI-2 Impact of the Camera sensor’s ‘HFOV’ on Trajectory Prediction
The derived RMSE values based on the camera sensor plotted against the prediction horizon for the three scenario categories are shown in Figs. 8, 9 and 10.
Figure 8: Impact on RMSE - Changing Camera’s HFOV: RLC Vs LLC - In front of Ego (US101)
Figure 9: Impact on RMSE - Changing Camera’s HFOV: RLC Vs LLC - In front of Lead (US101)
Figure 10: Impact on RMSE - Changing Camera’s HFOV: RLC Vs LLC - In front of Lead (Straight Road)
@@
In comparison to the observations made with the radar-based data, the gap between the RMSEs from the camera-based data and the ground truth RMSEs is much larger. It suggests that in general radar is likely to be a better-suited sensor modality for surrounding vehicle trajectory prediction in comparison to the camera. However, it is acknowledged that much more extensive data is required to solidify such a presumption.
@@
The camera sensor is known to struggle capturing rapid or complex movements accurately and this is likely the case for the target car and its surrounding vehicles during lane changes, especially with the cut-in manoeuvres included in the scenarios. The camera is also highly susceptible to occlusions. Figs. B.1, B.2, and B.3 in Supplementary Data B illustrate the detection of the target and the relevant surrounding vehicles between the two sensor modalities. Fig. B.2 shows how the occlusions erroneously elongate the duration of the lane change with the camera sensor, whereas, Fig. B.3 shows for the straight road scenarios, the algorithm even fails to identify the end of the lane change manoeuvres due to the camera failing to detect the target car following the lane change due to being occluded by the lead vehicle.
@@
For the lane changes to the right in scenarios ‘In front of the ego’ and ‘In front of the lead’, a HFOV setting of ${120}^{\circ}$ seems to provide the closest RMSEs to the ground truth RMSEs. However, for the left lane changes (LLCs) in those scenarios, the HFOV settings do not show a recognisable trend in terms of producing the RMSEs closest to the ground truth equivalents. The gap between the two RMSEs also remains considerably high towards the latter part of the prediction horizon.
@@
With the ‘StraightRd’ scenarios, for the right lane changes (RLCs), a HFOV of ${180}^{\circ}$ seem to produce the RMSEs closest to the base RMSEs whereas, for the LLCs, the closest to the base RMSEs is produced with a HFOV of ${120}^{\circ}$.
@@
However, with the two scenarios based on the US101 road segment, for the RLCs, a ${120}^{\circ}$ HFOV seems to work the best capturing all the relevant data without losing considerable resolution. However, the LLCs are inclined to be complex since they are manoeuvres into a high-speed lane and lack identifiable correlations with the HFOV changes. The considerably high RMSE gap is likely to be due to the camera’s susceptibility to dynamic scene changes and occlusions. The discrepancy in the optimal HFOV settings between the straight road and the US101-based scenarios is likely due to the differences in the road geometry.
@@
The common observation between the RLCs and the LLCs for all three scenario categories with the camera sensor data is that the HFOV of ${30}^{\circ}$ produces the largest gap between the two RMSE curves and the variation of the RMSE curves amongst the other HFOV settings not having such a considerable impact. I.e. for the camera sensor, varying FOV does not appear to impact the trajectory prediction model performance to a noticeable extent. The likely explanation for the above is the camera’s inability to capture critical inputs at the lowest HFOV of ${30}^{\circ}$, i.e. the largest gap between the RMSEs based on ground truths and the camera data is with a ${30}^{\circ}$ HFOV setting.
@@
Camera seems to require a broader view to compensate for the lack of visual data, especially with the lateral movements of traffic that happen during a lane change manoeuvre. On the other hand, lack of impact when increasing the camera’s HFOV beyond ${60}^{\circ}$ alludes to its ability to detect and process the relevant traffic objects quite robustly beyond a certain minimal HFOV such that further increases result in redundant data.
@@
Overall, the noticed difference in performance of the trajectory prediction model when the input data is sourced from two different sensor modalities is likely due to the fundamental operational differences between the two modalities. It is believed that the divergent strengths of the two sensor modalities, such as the radar’s ability to focus on the relevant traffic objects with a narrow HFOV whereas the camera’s dependency on visual data demanding a broader view to capture the required data, causes these observations. In addition, the high sensitivity of the camera to occlusions compared to the radar’s robustness to such environmental factors is a likely factor contributing to the difference. Hence, for an AAD vehicle deploying a trajectory prediction model such as the CS-LSTM, the inherent characteristics of the radar yield itself as an essential constituent of the perception sensor suite.
@@
Another interesting observation is that with the radar sensor inputs the prediction model seems to perform better for the RLC manoeuvres in comparison to the LLC manoeuvres. By comparing the radar’s detections of the target vehicle and its surrounding vehicles between the bi-directional lane changes as shown in Fig. C.1 in Supplementary Data C, it is seen that the target car and the surround car, NIO are detected less consistently in the LLC scenarios compared to the RLC scenarios. This explains the observation of slightly better prediction performances for the RLC manoeuvres since longer segments of missing data with the LLC manoeuvres contribute to poor predictions.
@@
As a crude litmus test to check the influence of missing data on the prediction model’s performance, data imputation is applied to the original radar data using linear interpolation. It is noted that with compensated data, regardless of the direction of lane change, the corresponding RMSEs along the prediction horizon are compatible with each other as shown in Supplementary Data D Figs. D.1 and D.2.
@@
In all scenarios studied, the RMSEs improve following data imputation, which makes for the case that the performance of the CS-LSTM degrades considerably with impaired detection of the target car and its surrounding vehicles. This is understandable since the loss of detection, even for a short duration, directly impacts the determination of the track history, which is the main input used by the DL model for making predictions. This highlights the need for further research to explore the robustness of the SOTA DL trajectory prediction models when deployed to operate using incomplete real-world data.
@@ VII Conclusion
Insights on how different types of perception sensors and their parameter configurations contribute to optimising the perception sensor suite of an AAD vehicle are highly desirable as the automotive industry progresses towards full automation. However, there are hardly any studies to understand the correlations between different perception sensor modalities/ sensor parameter settings and the performance impact on a DL-based perception function such as trajectory prediction. Such knowledge is essential to optimise the overall perception system efficiently and cost-effectively. It is identified that the lack of studies is mainly due to the challenge of obtaining sensor data with varying parameter settings whilst the rest of the variables remain constant in a highly dynamic driving environment.
As a way forward, a framework is introduced which combines a virtual simulation environment with sensor models to generate synthetic data for different sensor modalities. The framework involves all stages of the perception pipeline and a data analysis stage, offering maximum flexibility with its modular approach. Therefore, it can be used to conduct a multitude of different sensitivity studies. It facilitates quantitative sensitivity analyses between sensor configurations and the performance of a perception function based on an applicable metric to the selected perception function. Overall, the framework facilitates optimising the perception sensor suite by tailoring the sensor configuration to the needs of the perception algorithms implemented on the vehicle.
The contribution of our framework is demonstrated by using it to investigate the optimum parameter setting for the HFOV, in addition to selecting between the sensor modalities radar and camera for the application. The results suggest that a radar sensor with a narrow HFOV is the most suited sensor configuration for the perception algorithm tested.
@@
    Making efficient design choices of the sensor suite at the early stages of the development cycle
@@
    Defining minimum safety requirements to be complied with to achieve system homologation
Future work will extend the variety of the scenarios and the sensor parameters being studied. It is also planned to explore the optimal perception sensor selections and parameter configurations for different environmental conditions such as adverse weather and different road geometries.

### v2

@@ V-B Model Selection
V-C Inputs and outputs
Inputs to the model are the track histories of the target vehicle and its surrounding vehicles, taken for the time frames from $t=t-t_{h}$ to $t=t$, where $t$ is the time when the prediction is made and $t_{h}$ is the start of the track history. Input features consist of the $x,\ y$ co-ordinates as detailed in [5]. For our analysis, the track histories are derived in two ways.
@@ V-C Inputs and outputs
Positional co-ordinates from both approaches are converted to the target vehicle’s frame of reference before being used as inputs to the model. The output of the model infers the conditional probability distribution $P\left(Y\middle|X\right)$. Therefore, the inferred future $x,\ y$ positions at each future predicted time-frame are the parameters of a bivariate Gaussian probability distribution, where $\mu^{t}$ is the mean vector and $\Sigma^{t}$ is the covariance matrix.
@@
V-D Scenario Generation
Six driving scenarios are created with a target vehicle performing a lane change (LC) manoeuvre on a highway, four representing a road segment from the US Highway 101 (replicating the real-world environment of the dataset used for training the CS-LSTM model) and the remainder a straight highway road segment. The scenarios cover the most influential positions for a LC by a target vehicle to impact the trajectory of the ego car, a LC into the same lane as the ego car, directly in front of it, and in front of the lead vehicle to the ego car.
LC is identified as one of the safety-critical manoeuvres causing a high number of road accidents on highways [9, 10]. Therefore, developing knowledge on the impact of perception sensor configuration on LC prediction performance is of high value.
Although some OEMs have started localising the perception sensor suite based on the country, the experiments were designed to explore bi-directional LCs since the objective is overall safe driving on highways including LCs in relation to overtaking, exiting the highways, and to accommodate offensive driving.
  • Sc-01: Scenario LC to the Right - In front of the Ego car (US101)
  • Sc-02: Scenario LC to the Left - In front of the Ego Car (US101)
  • Sc-03: Scenario LC to the Right - In front of the Lead Car (US101)
  • Sc-04: Scenario LC to the Left - In front of the Lead Car (US101)
  • Sc-05: Scenario LC to the Right – In front of the Lead Car (Straight Road)
  • Sc-06: Scenario LC to the Left – In front of the Lead Car (Straight Road)
Detailed descriptions of Sc-01 and Sc-02 as per [38] and simulation recordings to visualise the scenarios are available at https://github.com/asha-gamage/perception4prediction_1
@@ V-D Scenario Generation
The ego car is represented by a typical passenger vehicle type along with several different vehicle types representing the surrounding vehicles. High-fidelity sensor models for radar and stereo camera from CM are mounted on the ego vehicle positioned as shown in Fig. 4. The radar model uses signal-to-noise ratio for detecting the surrounding objects with consideration given to propagation and atmospheric losses whereas the camera model uses a bounding box of the object for calculating occlusions to determine whether an object is detected. The detection confidence is compared with a predefined threshold, which for this study is set to 0.6. Tables I and II list the default sensor specifications for radar and camera sensors except for the technical parameter being changed, horizontal FOV (HFOV). The range of the sensor is fixed at 100 m.
@@
V-E Design of Experiments
A test run is established using a predefined scenario with one sensor modality and the sensor’s HFOV set as a variable. It is then simulated iteratively, varying the assigned value to the variable, HFOV, as listed in Table III. Although sensor technologies are continuously improving and expanding their operational boundaries, the experiment limits the technical parameters in line with the commercially available models to date [2, 35]. The parameter variations are selected in step-wise increments compared to a continuous range to enable identifying the optimal sensor settings based on commercially available sensors.
To study the influence of a sensor parameter in isolation due to the ration ale in Sec. V-A, all parameters, except the one being evaluated, are kept fixed. Identical sets of simulation test runs are conducted for the radar and camera sensors.
V-F Evaluation Metric
The metric root mean squared error (RMSE) as defined in Eq. 12 is used to analyse the results. The RMSEs on the predicted trajectories from the model using both ground truth data and sensor data are plotted against the prediction time horizon to compare the impact of different sensor modalities and parameter settings.
VI Results Interpretation and Discussion
@@ VI-1 Impact of the Radar sensor’s ‘HFOV’ on Trajectory Prediction
Fig. 5 shows the derived RMSE values based on radar data plotted against the prediction horizon of 5 s for the six scenarios.
Sc-01 Vs Sc-02
Sc-03 Vs Sc-04
Sc-05 Vs Sc-06
Figure 5: Impact on RMSE - Changing Radar’s HFOV: RLC Vs LLC
It is observed that the lowest HFOV setting of ${30}^{\circ}$ has produced the RMSEs closest to the ground truth-based RMSEs in all six scenarios explored: (i) in front of the ego car on US101, (ii) in front of the lead car on US101 and (iii) in front of the lead car on the straight road. The intuitive expectation is for the sensor-based data obtained with a larger HFOV to perform better when used as the input to the prediction model. However, the results exemplify that the RMSEs of the predicted trajectories are optimal when the measurements are taken with the lowest HFOV of ${30}^{\circ}$.
Discussion:
It is believed since the dominant manoeuvre of all six scenarios is a LC into the ego car’s lane, the narrow focus given by a smaller HFOV of the radar sensor enables the DL model to make a better prediction by reducing the complexity and unnecessary noise. Since for a LC, the highest impact on the target vehicle is imposed by the vehicles in its closest proximity and directly ahead, a narrow HFOV improves focus by reducing the irrelevant data of unrelated surrounding traffic. Another likely rationale is that the data captured at ${30}^{\circ}$ HFOV closely representing the relevant area of interest around the target vehicle. The CS-LSTM model uses the BEV representation of the scene with a predefined boundary around the target car. Therefore, when the sensor data is based on a wider HFOV, the inputs may appear distorted reducing the model’s performance.
The above observations highlight the significance of determining the value addition by deploying high-specification perception sensors within the sensor suite of an ADS as mentioned in [22]. Although the expectation is for high-specification sensors leading to high-quality data, the results of this experiment indicate that the performance of DL models might not always positively correlate with data from high-specification sensors.
@@ VI-2 Impact of the Camera sensor’s ‘HFOV’ on Trajectory Prediction
The derived RMSE values based on camera data plotted against the prediction horizon for the six scenarios are shown in Fig. 6.
Sc-01 Vs Sc-02
Sc-03 Vs Sc-04
Sc-05 Vs Sc-06
Figure 6: Impact on RMSE - Changing Camera’s HFOV: RLC Vs LLC
@@
In comparison to the observations made with radar-based data, the gaps between the camera-based RMSEs and ground truth-based RMSEs are much larger. It suggests that in general, radar is likely to be a better-suited sensor modality for surrounding vehicle trajectory prediction in comparison to camera. While further extensive data is needed to fully validate this presumption, it provides a strong foundation for further exploration.
Radar Detections
Camera Detections
Figure 7: Traffic Detection - RLC on US101 Highway
Radar Detections
Camera Detections
Figure 8: Traffic Detection - RLC on Straight Road
Discussion:
@@
The camera sensor is known to struggle capturing rapid or complex movements accurately and this is likely the case for the target car and its surrounding vehicles during LCs, especially with the cut-in manoeuvres included in the scenarios. The camera is also highly susceptible to occlusions. Fig. 7 shows how the occlusions leading to missing data with camera sensors erroneously elongate the duration of the LC, whereas, Fig. 8 shows for the straight road scenarios, the algorithm even fails to identify the end of LC manoeuvres due to the camera failing to detect the target car following the LC as it is occluded by the lead vehicle. The complete set of figures illustrating the detection of the target and the relevant surrounding vehicles between the two sensor modalities is available under Supplementary Data B at GitHub repository: https://github.com/asha-gamage/perception4prediction_1
@@
For LCs to the right in scenarios ‘In front of Ego’ and ‘In front of the Lead’, a HFOV setting of ${120}^{\circ}$ seems to provide the closest RMSEs to the ground truth RMSEs. However, for left lane changes (LLCs) in those scenarios, the HFOV settings do not show a recognisable trend in terms of producing RMSEs closest to the ground truth equivalents. The gap between the two RMSEs also remains considerably high towards the latter part of the prediction horizon.
@@
With Sc-05 and Sc-06, for the right lane changes (RLCs), a HFOV of ${180}^{\circ}$ seems to produce the RMSEs closest to the base RMSEs whereas, for the LLCs, the closest to the base RMSEs is produced with a HFOV of ${120}^{\circ}$.
@@
However, with four scenarios based on the US101 road segment, for the RLCs, a ${120}^{\circ}$ HFOV seems to work the best capturing all the relevant data without losing considerable resolution. However, the LLCs are inclined to be complex since they are manoeuvres into a high-speed lane and lack identifiable correlations with the HFOV changes. The considerably high RMSE gap is probably attributable to the camera’s susceptibility to dynamic scene changes and occlusions. The discrepancy in the optimal HFOV settings between the straight road and the US101-based scenarios is likely due to the differences in the road geometry.
@@
A consistent observation across all scenarios with the camera sensor data, is that a HFOV of ${30}^{\circ}$ generates the largest gap between the ground truth RMSE curve and the sensor parameter-based RMSE curve. In contrast, variations in the RMSE curves for other HFOV settings have a relatively minor impact. I.e. for camera, varying HFOV does not appear to impact the trajectory prediction model performance to a noticeable extent. The likely explanation is camera’s inability to capture critical inputs at the lowest HFOV of ${30}^{\circ}$ causing the largest gap between the RMSEs curves.
@@
Camera seems to require a broader view to compensate for the lack of visual data, especially with the lateral movements of traffic that happen during a LC manoeuvre. On the other hand, lack of impact when increasing the camera’s HFOV beyond ${60}^{\circ}$ alludes to its ability to detect and process the relevant traffic objects quite robustly beyond a certain minimal HFOV, such that further increases result in redundant data.
@@
Overall, the noticed difference in performance of the trajectory prediction model when the input data is sourced from two different sensor modalities is likely due to the fundamental operational differences between the two modalities. It is believed that the divergent strengths of the two sensor modalities, such as radar’s ability to focus on the relevant traffic objects with a narrow HFOV, and camera’s dependency on visual data demanding a broader view to capture the required data, causes these observations. In addition, the high sensitivity of the camera to occlusions compared to the radar’s robustness to such environmental factors is a likely factor contributing to the difference. Hence, for an ADS deploying a trajectory prediction model such as the CS-LSTM, radar becomes an essential component of the perception sensor suite due to its inherent characteristics.
@@
Another interesting observation is that with radar sensor inputs, the prediction model seems to perform better for the RLC manoeuvres in comparison to the LLC manoeuvres. By comparing the radar’s detections of the target vehicle and its surrounding vehicles between the bi-directional LCs, as shown in Fig. C.1 in Supplementary Data C, it is seen that the target car and the surround car, NIO, are detected less consistently in the LLC scenarios compared to the RLC scenarios. This explains the slightly better prediction performance for RLC manoeuvres, as the longer segments of missing data associated with LLC manoeuvres tend to degrade prediction accuracy.
@@
As a crude litmus test to check the influence of missing data on the model’s prediction performance, data imputation is applied to the original radar data using linear interpolation. It is noted that with compensated data, regardless of the direction of LC, the corresponding RMSEs along the prediction horizon are compatible with each other as shown in Supplementary Data D in GitHub repository.
@@
In all scenarios studied, the RMSEs improve following data imputation, which makes for the case that the performance of the CS-LSTM degrades considerably with impaired detection of the target car and its surrounding vehicles. This is understandable, as even short-duration detection losses directly affect the track history, which serves as the primary input to the CS-LSTM model for making predictions. This highlights the need for further research to explore the robustness of the SOTA DL trajectory prediction models when deployed with incomplete real-world data.
@@ VII Conclusion
Insights on how different types of perception sensors and their parameter configurations contribute to optimise the perception sensor suite of an ADS are highly desirable as the automotive industry progresses towards full automation. However, studies to understand the correlations between different perception sensor modalities/ parameter settings and the performance impact on a DL-based perception functions are rare. Such knowledge is essential to optimise the overall perception system efficiently and cost-effectively.
As a way forward, an evaluation framework is introduced which involves all stages of the perception pipeline, offering maximum flexibility with its modular approach to conduct a multitude of sensitivity studies. It facilitates quantitative sensitivity analyses between different sensor configurations and the performance of a perception function. Overall, the framework facilitates optimising the perception sensor suite by tailoring the sensor configuration to the needs of the perception algorithms implemented on the vehicle.
The contribution of our framework is demonstrated by using it to investigate the optimum parameter setting for the HFOV, and to select between the sensor modalities radar and camera for the application. The results suggest that a radar sensor with a narrow HFOV is the most suited sensor configuration for the tested perception algorithm.
@@
    Making efficient design choices for the sensor suite at the early stages of the development cycle
@@
    Defining minimum safety requirements to be complied to achieve system homologation
Future work will extend the variety of the scenarios and sensor parameters being studied by integrating enhanced sensor models. It is also planned to explore the optimal perception sensor selections and parameter configurations for different environmental conditions such as adverse weather and challenging road geometries.